"plot all the complex numbers in the complex plane whose absolute value is 4. " : "plot all the complex numbers in the complex plane whose absolute value is 4. "
Answer by stanbon(18993)

|a+bi| = sqrt(a^2+b^2)
If sqrt(a^2+b^2) = 4
then a^2+b^2 = 16
then {(a,b)} are all the points on a circle of radius 4 with center at the origin.
--------
Cheers,
Stan H.